# Krylov Complexity as a Probe for Chaos

### v2

@@ IV Explicit examples
Here $\sigma^{z,y,z}$ are Pauli matrices and $J,g$ and $h$ are constants which define the model. By rescaling one can set $J=1$. Then the nature of the model, namely whether it is chaotic or integrable, is solely determined by the value of the constants $g$ and $h$. In particular, for $gh\neq 0$ the model is non-integrable. In order to perform our numerical computations we will set $h=0.5,\,g=-1.05$ for chaotic case Bañuls et al. (2011), and $h=0,\,g=-1.05$ to extract behaviour in the integrable regime.
@@
In the chaotic limit of this Ising model (43), these states have been considered in Bañuls et al. (2011) to explore the nature of thermalization, in particular, whether the thermalization is strong or weak [^9] By making use of certain measures the nature of thermalization for different spin models has been studied in Sun et al. (2021a); Lin and Motrunich (2017); Chen et al. (2021); Bhattacharjee et al. (2022b); Nandy et al. (2024c); dos Prazeres and de Oliveira (2024); Alishahiha and Vasli (2024); Menzler and Jha (2024).. In the strong thermalization, the expectation value of a typical operator relaxes to the thermal value very fast, while for weak thermalization it strongly oscillates around the thermal value, though its time average attains the thermal value. It has been shown that although the initial state $|Y+\rangle$ exhibits strong thermalization, the initial states $|Z+\rangle$ and $|X+\rangle$ do not (see also Sun et al. (2021b)).
@@
It is straightforward to compute the Lanczos coefficients for these initial states from which one can compute complexity numerically101010 Lanczos coefficients for the model under consideration have also been computed inNoh (2021); Español and Wisniacki (2023); Trigueros and Lin (2022); Scialchi et al. (2024); Bhattacharya et al. (2023, 2022) . by using (19). The numerical results for the chaotic ($h=0.5$) and the integrable ($h=0$) cases are depicted in figures 1 and 2, respectively.
@@
Figure 1: Time evolution of complexity for different initial states $|X+\rangle,|Y+\rangle$ and $|Z+\rangle$ for the case where the model is chaotic; $g=-1.05,h=0.5$. The dashed lines represent the infinite time average of complexity. The computation is performed for $N=11$.
@@
Figure 2: Time evolution of complexity for different initial states $|X+\rangle,|Y+\rangle$ and $|Z+\rangle$ for the case where the model is integrable; $g=-1.05,h=0.5$. The dashed lines represent the infinite time average of complexity. The computation is performed for $N=11$.
We observe that the numerical results are perfectly compatible with the analytical argument we presented in the previous section. Indeed, the plots show distinctive behaviour of complexity for chaotic and integrable models. While that the complexity saturates to its saturation value given by the infinite time average of complexity at the saturation time for chaotic, in the integrable case it reaches this value from below at large times.
It is worth noting once again that the actual value of the complexity saturation does not provide a measure to probe whether the model is chaotic or integrable. As evident from the results presented in figures 1 and 2, the infinite time average of complexity for the chaotic system is not significantly different from that of the integrable models111111see Rabinovici et al. (2022a, b); Scialchi et al. (2024); Ballar Trigueros and Lin (2022); Español and Wisniacki (2023) for related discussions.
We have also observed a significant fluctuation around the saturation value, though it does not affect the physical behaviour of the complexity. The size and the nature of fluctuation depend on the initial states. Of course, the fluctuations become milder as we move to higher $N$.
Another interesting observation we have made is that although in the chaotic case and for the initial state $|Y+\rangle$ the complexity exhibits a clear peck before the saturation, it is not the case for other initial states. This is in line with the discussion in the previous section where we argued that for a generic initial state, complexity may or may not exhibit the peak. Even in cases where such a peak exists, it may not be pronounced enough. Of course, having a peak is guaranteed for particular initial states such as the thermofield double state as we have shown in the previous section. The second example we showed numerically is the initial state $|Y+\rangle$ in the chaotic regime of the Ising model (43).
@@
In order to crosscheck our analytic claim regarding the TFD state, we have also computed complexity for the initial state (41) for $\beta=0$. The result is shown in figure (3). From this figure one can immediately observe that the complexity for chaotic case does exhibit a peak before saturation. Furthermore, the behaviour of complexity for the integrable case also confirms the other part of our claim, namely it approaches the infinite time average from below and reaches it at late times.
@@
Figure 3: Time evolution of complexity for initial state given by (41) with $\beta=0$. The dashed lines represent the infinite time average of complexity. One observes that in the chaotic case the complexity exhibits a clear peak before saturation. The computation is performed for $N=11$.
@@
Thus far, our numerical computations have been centred on two specific points within the parameter space of the model (43), specifically $(g=-1.05,h=0)$ and $(g=-1.05,h=0.5)$. These points correspond to integrable and nearly maximally chaotic scenarios, leading to Poisson and Wigner distributions for level spacing, respectively. It would be more interesting to explore additional points and calculate the Krylov complexity, which would allow us to determine whether the system tends towards chaotic or integrable behaviours depending on the parametric closeness of the system to the aforementioned points. In pursuit of this, we have also evaluated the Krylov complexity for the initial state $|Y+\rangle$ at different values for $h$, yielding the results depicted in figure (4). Analysis of these figures reveals that as we progress towards higher values of $h$, the system moves away from chaos and leans towards integrability, as expected. This observation further supports our earlier conclusion in the preceding section that the $\sin$-term and level spacing play crucial roles in determining the behaviour of Krylov complexity at late times.
@@ V Discussions
Moreover, there are small fluctuations around the saturation value whose shapes also depend on the initial state. This confirms the postulates of Susskind (2016); Brown et al. (2017) for complexity saturation.
@@
It is important to highlight that achieving the desired saturation behaviour requires the function $A(E_{j},E_{k})$ to be smooth and not develop poles, a condition that holds for finite $n$. In infinite-dimensional systems, where $n$ can become very large, we can analyze the late-time behaviour of complexity by adopting the continuum limit. This approach enables us to compute the function $A(E_{j},E_{k})$, which often reveals a double-pole structure, resulting in linear growth of complexity at late times Alishahiha (2023); Alishahiha and Banerjee (2023).
Finally, it is worth noting that for a thermal initial state where $|c_{j}|^{2}=e^{-\beta E_{j}}/Z[\beta]$, the Krylov complexity can be expressed in a form which is reminiscent of the spectral complexity introduced in Iliesiu et al. (2022) (see also Camargo et al. (2024b)). Therefore, we would expect that the spectral complexity exhibits the same late-time behaviour as that of the Krylov complexity.

### v3

@@ IV Explicit examples
Here $\sigma^{z,y,z}$ are Pauli matrices and $J,g$ and $h$ are constants which define the model. By rescaling one can set $J=1$. Then the nature of the model, namely whether it is chaotic or integrable, is solely determined by the value of the constants $g$ and $h$. In particular, for $gh\neq 0$ the model is non-integrable. In order to perform our numerical computations we will set h=0.5,g=1.05h=0.5,\,g=-1.05 for chaotic case Banuls:2011vuw , and h=0,g=1.05h=0,\,g=-1.05 to extract behaviour in the integrable regime.
@@
In the chaotic limit of this Ising model (44), these states have been considered in Banuls:2011vuw to explore the nature of thermalization, in particular, whether the thermalization is strong or weak [^11] By making use of certain measures the nature of thermalization for different spin models has been studied in Sun:2020ybj ; Lin_2017 ; Chen:2021 ; Bhattacharjee:2022qjw ; Nandy:2023brt ; prazeres2024continuoustransitionweakstrong ; Alishahiha:2024rwm ; Menzler:2024atb .. In the strong thermalization, the expectation value of a typical operator relaxes to the thermal value very fast, while for weak thermalization it strongly oscillates around the thermal value, though its time average attains the thermal value. It has been shown that although the initial state $|Y+\rangle$ exhibits strong thermalization, the initial states $|Z+\rangle$ and $|X+\rangle$ do not (see also Sun:2020ybj ).
@@
It is straightforward to compute the Lanczos coefficients for these initial states from which one can compute complexity numerically121212 Lanczos coefficients for the model under consideration have also been computed inNoh_2021 ; Espa_ol_2023 ; Ballar_Trigueros_2022 ; Scialchi:2023bmw ; Bhattacharya:2023zqt ; Bhattacharya:2022gbz . by using (19). The numerical results for the chaotic ($h=0.5$) and the integrable ($h=0$) cases are depicted in figures 1131313 We normalized the complexity by its infinite-time average value $(C_{k}^{\infty})$ to better illustrate how complexity approaches this average in both integrable and chaotic cases..
@@
Figure 1: Time evolution of complexity for different initial states $|X+\rangle,|Y+\rangle$ and $|Z+\rangle$ for the case where the model is integrable ($h=0$) and chaotic ($h=0.5$). The computation is performed for $N=11$.
We observe that, although the numerical results are qualitatively consistent with the analytical argument presented in the previous section, different initial states can lead to significant fluctuations, making it difficult to clearly distinguish between chaotic and integrable behaviors. In particular, for the chaotic case and the initial state $|Y+\rangle$, the complexity shows a distinct peak before saturation, similar to that of the thermofield double state.
The magnitude and nature of these fluctuations depend on the initial states. Although the significant fluctuations around the saturation value do not affect the overall physical behavior of complexity, it is interesting to focus on initial states with milder fluctuations to observe clearer behavior, such as the $|Y+\rangle$ state in chaotic cases.
To identify an initial state with minimal fluctuations, we note that the inverse participation ratio (IPR), defined by
$\lambda^{-1}=\sum_{\alpha=1}^{\mathcal{D}}|c_{\alpha}|^{4},\;\;\;\;\;{\rm with}\;1\leq\lambda\leq{\mathcal{D}}\,$, (45)
may serve as a measure of the magnitude of fluctuations around the saturation value of complexity. Indeed, we have observed that the larger the IPR, the milder the fluctuations in complexity.
More precisely, when the IPR value is sufficiently large (for our model, when $\lambda>0.15\,\mathcal{D}$), fluctuations in complexity are suppressed, allowing for a clearer observation of its expected behavior in both chaotic and integrable states.
In Figure 2, we consider a special initial state where the IPR is large enough in both integrable and chaotic cases, ensuring that fluctuations do not obscure the complexity dynamics. The corresponding initial state is
$|\psi_{0}\rangle=|Y+\rangle_{1}\otimes|Z-\rangle_{2}\otimes|Y+\rangle_{3}\otimes|Z-\rangle_{4}\otimes\cdots\,.$ (46)
@@
Figure 2: Time evolution of complexity for initial state given by (46) for the case where the model is integrable ($h=0$) and chaotic ($h=0.5$). The computation is performed for $N=11$.
@@
In order to crosscheck our analytic claim regarding the TFD state, we have also computed complexity for the initial state (42) for $\beta=0$ [^14] It is worth noting that this initial state has maximal IPR (i.e., equal to $\cal D$}), making it a very suitable choice for studying the behavior of complexity saturation.. The result is shown in figure (3). From this figure one can immediately observe that the complexity for chaotic case does exhibit a peak before saturation. Furthermore, the behavior of complexity for the integrable case also confirms the other part of our claim, namely it approaches the infinite time average from below and reaches it at late times.
@@
Figure 3: Time evolution of complexity for initial state given by (42) with $\beta=0$. One observes that in the chaotic case the complexity exhibits a clear peak before saturation. The computation is performed for $N=11$.
These plots show distinctive behavior of complexity for chaotic and integrable models. While the complexity saturates to its saturation value given by the infinite time average of complexity at the saturation time for chaotic, in the integrable case it reaches this value from below at large times.
@@
Thus far, our numerical computations have been centred on two specific points within the parameter space of the model (44), specifically (g=1.05,h=0)(g=-1.05,h=0) and (g=1.05,h=0.5)(g=-1.05,h=0.5). These points correspond to integrable and nearly maximally chaotic scenarios, leading to Poisson and Wigner distributions for level spacing, respectively. It would be more interesting to explore additional points and calculate the Krylov complexity, which would allow us to determine whether the system tends towards chaotic or integrable behaviours depending on the parametric closeness of the system to the aforementioned points. In pursuit of this, we have also evaluated the Krylov complexity for the initial state $|Y+\rangle$ at different values for $h$, yielding the results depicted in figure (4). Analysis of these figures reveals that as we progress towards higher values of $h$, the system moves away from chaos and leans towards integrability, as expected. This observation further supports our earlier conclusion in the preceding section that the $\sin$-term and level spacing play crucial roles in determining the behaviour of Krylov complexity at late times.
@@ V Discussions
Indeed, we established a connection between complexity and level spacing statistics, a well-known measure of chaos. This relationship enabled us to explicitly demonstrate why and how complexity can serve as an effective indicator of chaos.
@@
Moreover, there are small fluctuations around the saturation value whose shapes also depend on the initial state. This confirms the postulates of Susskind:2014moa ; Brown:2016wib for complexity saturation. We also found that these fluctuations are directly related to the inverse participation ratio (IPR). A state with a sufficiently large IPR exhibits the expected behavior more clearly and reliably.
@@
It is important to highlight that achieving the desired saturation behaviour requires the function $A(E_{j},E_{k})$ to be smooth and not develop poles, a condition that holds for finite $n$. In infinite-dimensional systems, where $n$ can become very large, we can analyze the late-time behaviour of complexity by adopting the continuum limit. This approach enables us to compute the function $A(E_{j},E_{k})$, which often reveals a double-pole structure, resulting in linear growth of complexity at late times Alishahiha:2022nhe ; Alishahiha:2022anw .
Finally, it is worth noting that for a thermal initial state where $|c_{j}|^{2}=e^{-\beta E_{j}}/Z[\beta]$, the Krylov complexity can be expressed in a form which is reminiscent of the spectral complexity introduced in Iliesiu:2021ari (see also Camargo:2023eev ). Therefore, we would expect that the spectral complexity exhibits the same late-time behaviour as that of the Krylov complexity.